# Collapsing regular Riemannian foliations with flat leaves

### v1

@@ 2.5. Alexandrov spaces
We now recall the definition of an Alexandrov space. A locally compact and locally complete inner metric space $(X,d_{X})$, is called a local Alexandrov space if we have the following: For any $x\in X$ there exist $\kappa_{x}\in\mathbb{R}$ and an open neighborhood $U\subset X$ of $x$, such that for each geodesic triangle $\triangle$ contained in $U$, there exists a geodesic triangle $\widetilde{\triangle}$ in $S^{2}_{\kappa_{x}}$, the $2$-dimensional surface of constant sectional curvature $\kappa_{x}$, with edges having the same lengths as the edges of $\triangle$ satisfying the following condition: For a vertex $y\in\triangle$, $\tilde{y}$ the corresponding vertex in $\widetilde{\triangle}$, and $w$ any point in the opposing edge to $y$ in $\triangle$ with corresponding point $\tilde{w}$ in the opposing edge to $\tilde{y}$ in $\widetilde{\triangle}$, then
$d(y,w)\geq d_{S^{2}_{\kappa_{x}}}(\tilde{y},\tilde{w}).$
We call $x$ a point of $X$ with curvature at least $\kappa_{x}$. When there exist $\kappa\in\mathbb{R}$ such that any point of $X$ is a point with curvature at least $\kappa$, we say that $(X,d_{X})$ is an Alexandrov space of curvature at least $\kappa$, and we denote it by $\operatorname{curv}(X)\geq\kappa$. We denote the collection of all Alexandrov spaces with curvature at least $\kappa\in\mathbb{R}$ by $\mathrm{Alex}(\kappa)$.
@@
þ Let $(X,d_{X})$ be a compact $n$-dimensional Alexandrov space of $\operatorname{curv}(X)\geq\kappa$. Then there exists an $\epsilon=\epsilon(X)>0$ such that for any $n$-dimensional Alexandrov space $(Y,d_{Y})$ of $\operatorname{curv}(Y)\geq\kappa$ with $d_{GH}(X,Y)<\epsilon$, $Y$ is homeomorphic to $X$.
@@
þ The homeomorphism in þ2.6 can be taken so that it preserves the stratification of $X$ and $Y$ by extremal subsets (see [32], [30, Theorem 4.3]).
@@ 2.6. Equivariant convergence
We denote by $\mathcal{M}$ the set of all isometry classes of pointed metric spaces $(X,d_{X},x)$ such that for each $r>0$ the open ball $B_{r}(x)$ is relatively compact. We define the following notion of convergence on $\mathcal{M}$ as follows:
@@
We say that a a pointed sequence $\{(X_{i},d_{X_{i}},x_{i})\}_{i\in\mathbb{N}}\subset\mathcal{M}$ converges to $(Y,d_{Y},y)\in\mathcal{M}$ if for each $r>0$ and each $\varepsilon>0$ there exists $N\in\mathbb{N}$, such that for $n\geq N$ there exists a map $f_{n}\colon B_{r}(x_{n})\to X$ such that
@@
þ When we consider a sequence of compact spaces $(X_{i},d_{X_{i}})$ and $(Y,d_{Y})$, then convergence with respect to the Gromov-Hausdorff distance is equivalent to convergence in the pointed Gromov-Hausdorff sense.
Moreover, for compact spaces we have a nice characterization of converge in the Gromov-Hausdorff sense. Recall that given $\varepsilon>0$, a subset $S\subset X$ of a metric space is an $\varepsilon$-net if we have that $d_{X}(x,S)=\inf\{d_{X}(x,s)\mid s\in S\}\leq\varepsilon$. Given $X$, $Y$ two compact metric spaces, and $\varepsilon,\delta>0$, we say that $X$ and $Y$ are $(\varepsilon,\delta)$-approximations of each other if there exists $\{x_{i}\}_{i=1}^{N}\subset X$, $\{y_{i}\}_{i=1}^{N}\subset Y$ such that
@@
When $\varepsilon=\delta$, we say that $X$ and $Y$ are $\varepsilon$-approximations of each other. With this we can write the characterization of convergence in the Gromov-Hausdorff sense for compact spaces:
Proposition 2.10 (Proposition 7.4.11).
þ Let $X$ and $Y$ be compact metric spaces. Then we have
@@
We define the following topology on $\mathcal{M}_{\mathrm{eq}}$: Given $(X,d_{X},H,x)$, $(Y,d_{Y},H^{\prime},y)$ in $\mathcal{M}_{\mathrm{eq}}$ and $\varepsilon>0$, we define an $\varepsilon$-equivariant pointed Gromov-Hausdorff approximation between $X$ and $Y$ to be a triple of functions $(f,\phi,\psi)$, {linenomath}
@@
þ Let $\{(X_{i},d_{X_{i}},H_{i},x_{i})\}_{i\in\mathbb{N}}\subset\mathcal{M}_{\mathrm{eq}}$ have limit $(Y,d_{Y},H^{\prime},y)\in\mathcal{M}_{\mathrm{eq}}$ with respect to the equivariant pointed Gromov-Hausdorff distance. Then we have that $(Y/H^{\prime},d_{Y}^{\ast},y^{\ast})$ is the limit of the sequence $\{(X_{i}/H_{i},d_{X_{i}}^{\ast},x_{i}^{\ast})\}_{i\in\mathbb{N}}$ in the sense of pointed Gromov-Hausdorff convergence.
@@
þ Assume that for a sequence $\{(X_{i},d_{X_{i}},H_{i},x_{i})\}_{i\in\mathbb{N}}$ in $\mathcal{M}_{\mathrm{eq}}$ we have that $(Y,d_{Y},y)\in\mathcal{M}$ is the limit of $(X_{i},d_{X_{i}},x_{i})$ in the pointed Gromov-Hausdorff sense. Then there exist $H^{\prime}\subset\mathrm{Isom}(Y,d_{Y})$ a closed subgroup acting effectively on $Y$, and a subsequence $\{(X_{i_{k}},d_{X_{i_{k}}},H_{i_{k}},x_{i_{k}})\}_{k\in\mathbb{N}}$ converging to $(Y,d_{Y},H^{\prime},y)$ with respect to the pointed equivariant Gromov-Hausdorff topology.
@@ 2.7. $\mathsf{N}$- and $\mathsf{F}$-structures
The concept of an $\mathsf{F}$-structure was introduced by Cheeger and Gromov in [6] generalizing the notion of torus actions on manifolds. Moreover they proved that for $\mathsf{F}$-structure of positive rank on a compact manifold we can find a family of Riemannian metrics $\{g_{t}\mid 1\geq t>0\}$ on the manifold such that the manifold collapses with bounded curvature and diameter to a metric space. That is, the injectivity radius of $(M,g_{t})$ goes to 00 as $t\to 0$, there exist $\lambda\leq\Lambda$ and $0<D$ such that $\lambda\leq\mathrm{Sec}(g_{t})\leq\Lambda$ and $\mathrm{Diam}(M,g_{t})\leq D$ for all $0<t\leq 1$.
@@
    The action of $\mathrm{O}(m)$ preserves the fibers and the structure group of the fibration.
A pure $\mathsf{N}$-structure induces a partition of $M$ into submanifolds called “orbits”. We refer to this decomposition of $M$ as a pure $\mathsf{N}$-structure on $M$. An $\mathsf{N}$-structure has positive rank if the orbits have positive dimension.
@@
þ An $\mathsf{N}$-structure $\tilde{\eta}\colon F(M)\to B_{\tilde{\eta}}$ on $M$ a manifold with finite fundamental group coincides with its canonical torus bundle $\tilde{f}\colon F(M)\to B_{\tilde{f}}$. That is, $B_{\tilde{\eta}}=B_{\tilde{f}}$ and $\tilde{\eta}=\tilde{f}$. Moreover, when $M$ is simply connected then an $\mathsf{N}$-structure on $M$ is given by a torus action on $M$.
@@
In [5] the authors showed that a sufficiently collapsed manifold $(M,g)$ admits an $\mathsf{N}$-structure. And moreover there exist an $\mathsf{N}$-invariant Riemannian metric on $M$ related to $g$.
@@
þ For $m>2$ and $D>0$ we denote by $\mathfrak{M}(m,D)$ the class of $m$-dimensional compact connected Riemannian manifolds $(M,g)$ with $|\mathrm{Sec}(g)|\leq 1$, and $\mathrm{Diam}(M,g)\leq D$.
@@
þ Let the assumptions be as in þ2.15. Then the nearby metric $g^{\varepsilon}$ can be chosen to satisfy in addition
$\min\sec(g)-\varepsilon\leq\sec(g^{\varepsilon})\leq\max\sec(g)+\varepsilon.$
In particular the metric $g^{\varepsilon}$ in þ2.15 induces an inner metric $d_{g^{\varepsilon}}^{\ast}$ on the quotient spaces of the $\mathsf{N}$-structures making them into Alexandrov spaces of curvature at least $\min\sec(g)-\varepsilon$.
Combining Theorems 2.6, 2.14, 2.15, and 2.16 we obtain the following result:
@@
þ Assume that $(M_{n},g_{n})$ is a sequence of simply-connected compact Riemannian $m$-manifolds with sectional curvature bounds $\lambda\leq\sec(g_{n})\leq\Lambda$ and diameters $\operatorname{diam}(g_{n})\leq D$ which collapses with bounded curvature. Then, given any $\varepsilon>0$, for $n$ sufficiently large the following holds:
@@
    There exists a smooth global effective $T^{k}$ action on $M_{n}$ with empty fixed-point set all of whose orbits have diameter less than $\varepsilon$ with respect to $g_{n}$.
@@
    $e^{-\varepsilon}g_{n}<g^{\varepsilon}_{n}<e^{\varepsilon}g_{n},\quad\lambda-\varepsilon\leq\sec(g^{\varepsilon}_{n})\leq\Lambda+\varepsilon.$
@@
    Moreover assume that $(M_{n},g_{n})$ converges to an Alexandrov space $X$ of dimension $(m-k)$ in the Gromov-Hausdorff topology. Then for the orbit space $M_{n}/T^{k}$ equipped with the metric induced by $g^{\varepsilon}_{n}$, we have that the Gromov-Hausdorff distance between $X$ and $M_{n}/T^{k}$ is less than $\varepsilon$.
@@ 3. Proofs of Main Theorems
In this section we give the proofs of some of our main results. We first present the proof of þB. We then present the proof of þC. With this we are able to give the proof of þA.
3.1. Proof of Theorem B
We start by giving a proof of þB.
Proof of þB.

### v2

@@ 2.5. Alexandrov spaces
We now recall the definition of an Alexandrov space. Let $(X,d_{X})$ be a locally compact and locally complete inner metric space. A point $x\in X$ is said to have curvature at least $\kappa$, for some $\kappa\in\mathbb{R}$, if there exist $U_{x}\subset X$ an open neighborhood of $x$ such that the following holds: Given any geodesic triangle $\Delta\subset U_{x}$ there exists a geodesic triangle $\widetilde{\triangle}$ in the $2$-dimensional surface of constant sectional curvature $\kappa$ denoted by $S^{2}_{\kappa}$, with edges having the same lengths as the edges of $\triangle$, and such that given $y\in\triangle$ a vertex and $w$ any point in the opposing edge to $y$ in $\triangle$, for $\tilde{y}$ and $\tilde{w}$ the unique corresponding points in $\widetilde{\triangle}$ to $y$ and $w$ respectively, then we have
[…]
A local Alexandrov space $(X,d_{X})$ is a locally compact and locally complete inner metric space such that for each point $x\in X$ there exist $\kappa_{x}\in\mathbb{R}$ so that $x$ has curvature at least $\kappa_{x}$.
In the case when there exists $\kappa\in\mathbb{R}$ such that any point of $X$ is a point with curvature at least $\kappa$, we say that $(X,d_{X})$ is an Alexandrov space of curvature at least $\kappa$, and we denote it by $\operatorname{curv}(X)\geq\kappa$. We denote the collection of all Alexandrov spaces with curvature at least $\kappa\in\mathbb{R}$ by $\mathrm{Alex}(\kappa)$.
@@
Let $(X,d_{X})$ be a compact $n$-dimensional Alexandrov space of $\operatorname{curv}(X)\geq\kappa$. Then there exists an $\epsilon=\epsilon(X)>0$ such that for any $n$-dimensional Alexandrov space $(Y,d_{Y})$ of $\operatorname{curv}(Y)\geq\kappa$ with $d_{GH}(X,Y)<\epsilon$, $Y$ is homeomorphic to $X$.
An extremal subset $E$ of an Alexandrov space $(X,d_{X})$ is a closed subset subset which is preserved under the gradient flow of $d_{X}(p,\cdot)$ for all $p\in X$ (see [32]).
Example 2.8.
Let $H$ act effectively smoothly by isometries on a compact Riemannian manifold $(M,g)$, consider $\pi\colon M\to M/H$ and equip $M/H$ with the induced metric $d^{\ast}$. Then $(M/H,d^{\ast})$ is an Alexandrov space. Let $\Gamma\subset H$ be a closed subgroup, and $F_{\Gamma}$ be the set of points fixed by $\Gamma$. Then by [42, 4.2 Proposition] $\pi(F_{\Gamma})$ is an extremal subset of $(M/H,d^{\ast})$.
@@
The homeomorphism in ?THM? LABEL:T:_Perelmans_stability can be taken so that it preserves the stratification of $X$ and $Y$ by extremal subsets (see [34],[31, Theorem 4.3]).
@@ 2.6. Equivariant convergence
We denote by $\mathcal{M}$ the set of all isometry classes of pointed metric spaces $(X,d_{X},x)$ such that for each $r>0$ the open ball $B_{r}(x)$ is relatively compact. We define the following notion of convergence on $\mathcal{M}$:
@@
We say that a a pointed sequence $\{(X_{i},d_{X_{i}},x_{i})\}_{i\in\mathbb{N}}\subset\mathcal{M}$ converges to $(Y,d_{Y},y)\in\mathcal{M}$ if for each $r>0$ and each $\varepsilon>0$ there exists $N\in\mathbb{N}$, such that for $n\geq N$ there exists a map $f_{n}\colon B_{r}(x_{n})\to X$ for which the following hold:
@@
When we consider a sequence of compact spaces $(X_{i},d_{X_{i}})$ and $(Y,d_{Y})$, then convergence with respect to the Gromov-Hausdorff distance is equivalent to convergence in the pointed Gromov-Hausdorff sense, in the following sense: in the case when $(X_{i},d_{X_{i}})$ converges in the Gromov-Hausdorff sense to $(Y,d_{Y})$, given $y\in Y$ there exists a sequence $\{x_{i}\mid x_{i}\in X_{i}\}_{i\in\mathbb{N}}$ such that $(X_{i},d_{X_{i}},x_{i})$ converges in the pointed-Gromov-Hausdorff sense to $(Y,d_{Y},y)$ (see [4, Exercise 8.12]).
Moreover, for compact spaces we have a nice characterization of convergence in the Gromov-Hausdorff sense. Recall that given $\varepsilon>0$, a subset $S\subset X$ of a metric space is an $\varepsilon$-net if we have that $d_{X}(x,S)=\inf\{d_{X}(x,s)\mid s\in S\}\leq\varepsilon$. Given $X$, $Y$ two compact metric spaces, and $\varepsilon,\delta>0$, we say that $X$ and $Y$ are $(\varepsilon,\delta)$-approximations of each other if there exists $\{x_{i}\}_{i=1}^{N}\subset X$, $\{y_{i}\}_{i=1}^{N}\subset Y$ such that
@@
When $\varepsilon=\delta$, we say that $X$ and $Y$ are $\varepsilon$-approximations of each other. With this we can write the characterization of convergence in the Gromov-Hausdorff sense for compact spaces.
Proposition 2.12 (Proposition 7.4.11 in [4]).
Let $X$ and $Y$ be compact metric spaces. Then we have
@@
We define the following topology on $\mathcal{M}_{\mathrm{eq}}$: Given $(X,d_{X},H,x)$, $(Y,d_{Y},H^{\prime},y)$ in $\mathcal{M}_{\mathrm{eq}}$ and $\varepsilon>0$, we define an $\varepsilon$-equivariant pointed Gromov-Hausdorff approximation between $X$ and $Y$ to be a triple of functions $(f,\phi,\psi)$,
@@
Let $\{(X_{i},d_{X_{i}},H_{i},x_{i})\}_{i\in\mathbb{N}}\subset\mathcal{M}_{\mathrm{eq}}$ have limit $(Y,d_{Y},H^{\prime},y)\in\mathcal{M}_{\mathrm{eq}}$ with respect to the equivariant pointed Gromov-Hausdorff distance. Then we have that $(Y/H^{\prime},d_{Y}^{\ast},y^{\ast})$ is the limit of the sequence $\{(X_{i}/H_{i},d_{X_{i}}^{\ast},x_{i}^{\ast})\}_{i\in\mathbb{N}}$ in the sense of pointed Gromov-Hausdorff convergence.
@@
Assume that for a sequence $\{(X_{i},d_{X_{i}},H_{i},x_{i})\}_{i\in\mathbb{N}}$ in $\mathcal{M}_{\mathrm{eq}}$ we have that $(Y,d_{Y},y)\in\mathcal{M}$ is the limit of $(X_{i},d_{X_{i}},x_{i})$ in the pointed Gromov-Hausdorff sense. Then there exist $H^{\prime}\subset\mathrm{Isom}(Y,d_{Y})$ a closed subgroup acting effectively on $Y$, and a subsequence $\{(X_{i_{k}},d_{X_{i_{k}}},H_{i_{k}},x_{i_{k}})\}_{k\in\mathbb{N}}$ converging to $(Y,d_{Y},H^{\prime},y)$ with respect to the pointed equivariant Gromov-Hausdorff topology.
@@ 2.7. $\mathsf{N}$- and $\mathsf{F}$-structures
The concept of an $\mathsf{F}$-structure (already defined in Section 1) was introduced by Cheeger and Gromov in [6] generalizing the notion of torus actions on manifolds. Moreover they proved that for given $\mathsf{F}$-structure whose orbits have positive dimension on a compact manifold we can find a family of Riemannian metrics $\{g_{t}\mid 1\geq t>0\}$ on the manifold such that the manifold collapses with bounded curvature. That is, the injectivity radius of $(M,g_{t})$ goes to 0 as $t\to 0$, and there exist $\lambda\leq\Lambda$ such that $\lambda\leq\mathrm{Sec}(g_{t})\leq\Lambda$ for all $0<t\leq 1$.
@@
    The action of $\mathrm{O}(m)$ preserves the fibers and the structure group of the fibration $\tilde{\eta}$.
A pure $\mathsf{N}$-structure induces a partition $\mathscr{O}_{\tilde{\eta}}$ of $M$ into submanifolds called orbits which are the images of the $\tilde{\eta}-$fibers under the projection $\pi\colon F(M)\to M$ (observe that the “orbits” of $\mathscr{O}$ are not orbits of a group action, but they are defined by orbits of the action of $N$ on $F(M)$). We refer to this decomposition of $M$ as a pure $\mathsf{N}$-structure on $M$. An $\mathsf{N}$-structure has positive rank if the orbits have positive dimension.
@@
An $\mathsf{N}$-structure $\tilde{\eta}\colon F(M)\to B_{\tilde{\eta}}$ on $M$ a manifold with finite fundamental group coincides with its canonical torus bundle $\tilde{f}\colon F(M)\to B_{\tilde{f}}$. That is, $B_{\tilde{\eta}}=B_{\tilde{f}}$ and $\tilde{\eta}=\tilde{f}$. Moreover, when $M$ is simply connected then an $\mathsf{N}$-structure on $M$ is given by a torus action on $M$.
@@
In [5] the authors showed that a sufficiently collapsed manifold $(M,g)$ admits an $\mathsf{N}$-structure. And moreover there exists an $\mathsf{N}$-invariant Riemannian metric on $M$ related to $g$.
@@
For $m>2$ and $D>0$ we denote by $\mathfrak{M}(m,D)$ the class of $m$-dimensional compact connected Riemannian manifolds $(M,g)$ with $|\mathrm{Sec}(g)|\leq 1$, and $\operatorname{Diam}(M,g)\leq D$.
@@
Let the assumptions be as in ?THM? LABEL:T:_Existence_of_N_structures_on. Then the nearby metric $g^{\varepsilon}$ can be chosen to satisfy in addition
$\min\mathrm{Sec}(g)-\varepsilon\leq\mathrm{Sec}(g^{\varepsilon})\leq\max\mathrm{Sec}(g)+\varepsilon.$
In particular the metric $g^{\varepsilon}$ in ?THM? LABEL:T:_Existence_of_N_structures_on induces an inner metric $d_{g^{\varepsilon}}^{\ast}$ on the quotient spaces $M/\mathscr{O}_{\tilde{\eta}_{\varepsilon}}$ of the $\mathsf{N}$-structures making them into Alexandrov spaces of curvature at least $\min\mathrm{Sec}(g)-\varepsilon$.
Example 2.19.
Given an effective smooth $T^{k}$-action by isometries on a Riemannian manifold $(M,g)$, let $\Gamma\subset T^{k}$ be an isotropy group of some point $p\in M$. The orbit type stratum $\Sigma_{\Gamma}^{\ast}\subset M/T^{k}$ correspond to all the orbits in $M/T^{k}$ that have isotropy $\Gamma$. Observe that the closure of $\Sigma_{\Gamma}^{\ast}$ corresponds to the set of all orbits whose isotropy group contains $\Gamma$ (See ?THM? LABEL:Ex:_Orbit_type_stratification).
Combining ?THM? LABEL:Ex:_orbit_type_stratification_torus_actions, Theorems LABEL:T:_Perelmans_stability, LABEL:T:_N-structure_over_manifold_with_finite_fundamental_group_are_given_by_T-structure, LABEL:T:_Existence_of_N_structures_on, and LABEL:T:_Rong_bounds_sectional_curvature we obtain the following result:
@@
Assume that $(M_{n},g_{n})$ is a sequence of simply-connected compact Riemannian $m$-manifolds with sectional curvature bounds $\lambda\leq\mathrm{Sec}(g_{n})\leq\Lambda$ and diameters $\operatorname{Diam}(g_{n})\leq D$ which collapses with bounded curvature. Then, given any $\varepsilon>0$, for $n$ sufficiently large the following holds:
@@
    There exists a smooth global effective $T^{k}$ action on $M_{n}$ with empty fixed-point set, all of whose orbits have diameter less than $\varepsilon$ with respect to $g_{n}$.
@@
    $e^{-\varepsilon}g_{n}<g^{\varepsilon}_{n}<e^{\varepsilon}g_{n},\quad\lambda-\varepsilon\leq\mathrm{Sec}(g^{\varepsilon}_{n})\leq\Lambda+\varepsilon.$
@@
    Moreover assume that $(M_{n},g_{n})$ converges to an Alexandrov space $X$ of dimension (mk)(m-k) in the Gromov-Hausdorff topology. Then for the orbit space $M_{n}/T^{k}$ equipped with the metric induced by $g^{\varepsilon}_{n}$, we have that the Gromov-Hausdorff distance between $X$ and $M_{n}/T^{k}$ is less than $\varepsilon$.
@@ 3. Proofs of Main Theorems
In this section we give the proofs of some of our main results. We first present the proof of ?THM? LABEL:MT:_Collapse_of_flat_foliations. We then present the proof of ?THM? LABEL:MT:_Approximating_flat_foliations_by_N-structures. With this we are able to give the proof of ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions.
3.1. Proof of Theorem LABEL:MT:_Collapse_of_flat_foliations
We start by giving a proof of ?THM? LABEL:MT:_Collapse_of_flat_foliations. The strategy of the proof is to explicitly define the sequence of collapsing metrics, and then verify that they collapse with bounded curvature and diameter.
Proof of ?THM? LABEL:MT:_Collapse_of_flat_foliations.